OpenStudy (anonymous):

Hi again, I have another "word" problem that is confusing me. If someone would be willing to help I would appreciate it:) Here is the question: On Oct. 12 a state income tax payment was made, which was due on June 30. The penalty was 8% simple interest on the unpaid tax of $909. Find the penalty using exact interest. Thanks in advance for any help you guys:) denise

OpenStudy (anonymous):

I think the question is poorly written but that is typical homework here. I assumed I needed to figure out the number of days and divide those by 365, then multiply by .08. Then by 909. I keep getting 52 something or other and the answer is supposed to be 20.72. $909(.08*261/365)

OpenStudy (campbell_st):

isn't you time period July 31 days August 31 days September 30 days October 12 days so the time is n = 104/365 so I = 909 * 0.08 * 104/365 just a thought.

OpenStudy (anonymous):

Ok, I was using this chart and it is way confusing. I will do these the way you did with the numbers of days in the months, thank you much, and I will see if I get the right answer;) denise

OpenStudy (anonymous):

Exactly right:) Thank you Campbell, denise